a spice mixture is 25% thyme.How many grams of thyme must be added to a 12 g. of mixture to increase the thyme content to 40%?
Answer by stanbon(75887)

a spice mixture is 25% thyme.How many grams of thyme must be added to a 12 g. of mixture to increase the thyme content to 40%?
---------------
Equation:
thyme + thyme = thyme
0.25x + 1.0(12-x) = 0.40*12
---
Multiply thru by 100 to get:
25x + 100*12-100x = 40*12
-75x = -60*12
x = 9.6 g (amt of 25% mixture needed in the final mix)
12-x = 2.4 g (amt of pure thyme needed in the final mix)
===============
Cheers,
Stan H.